OpenStudy (rissyroo118):

x y 1 320 2 340 3 360 4 380 5 400 A. Write an equation that represents the table shown.

OpenStudy (rissyroo118):

help me pleaseee!

geerky42 (geerky42):

Do you know what y is at x = 0?

OpenStudy (rissyroo118):

no i do not

jhonyy9 (jhonyy9):

so from these result for x=0 y=300 x_1 =1 --- y_1 = 300 + 1*20 = 320 x_2 =2 --- y_2 = 300 +2*20 =340 - - - x_n = n ---- y_n = 300 +n*20 hope this is understandably easy